### v1

# Nested subspace learning with flags

Abstract
Many machine learning methods look for low-dimensional representations of the data. The underlying subspace can be estimated by first choosing a dimension $q$ and then optimizing a certain objective function over the space of $q$-dimensional subspaces—the Grassmannian. Trying different $q$ yields in general non-nested subspaces, which raises an important issue of consistency between the data representations. In this paper, we propose a simple trick to enforce nestedness in subspace learning methods. It consists in lifting Grassmannian optimization problems to flag manifolds—the space of nested subspaces of increasing dimension—via nested projectors. We apply the flag trick to several classical machine learning methods and show that it successfully addresses the nestedness issue.
Keywords: subspace learning, Grassmann manifolds, flag manifolds, nested subspaces
1 Introduction
Finding low-dimensional representations of datasets is a quite common objective in machine learning, notably in virtue of the curse of dimensionality (Bellman, 1984). A classical way of building such a representation is by searching for a low-dimensional subspace that well represents the data, where “well” is defined by an application-dependent criterion (Cunningham and Ghahramani, 2015). Here is a non-exhaustive list of examples of subspace learning problems. Principal component analysis (PCA) (Jolliffe, 2002) searches for a low-dimensional subspace that minimizes the average squared Euclidean distance to the data. Robust subspace recovery (RSR) (Lerman and Maunu, 2018b) minimizes the average (absolute) Euclidean distance to the data, which is less sensitive to outliers. Many matrix decomposition methods, like robust PCA (Candès et al., 2011) and matrix completion (Keshavan et al., 2010; Candès and Recht, 2012) look for low-rank approximations of the data matrix, which can be decomposed into a product of subspace and coordinate matrices. Trace ratio (TR) (Ngo et al., 2012) refers to a wide class of problems that look for subspaces making a tradeoff between desired yet antagonist properties of low-dimensional embeddings—for instance Fisher’s linear discriminant analysis (LDA) (Fisher, 1936) which seeks to maximize the between-class variance while minimizing the within-class variance. Domain adaptation methods learn some domain-invariant subspaces (Baktashmotlagh et al., 2013) by minimizing the projected maximum mean discrepancy (MMD) (Gretton et al., 2012) between the source and target distributions. Subspace tracking methods incrementally minimize a distance between the current subspace and the available data (Balzano et al., 2010). Subspaces can also be estimated not from the data but from their adjacency matrix or graph Laplacian, as done in the celebrated Laplacian eigenmaps (Belkin and Niyogi, 2003) and spectral clustering (Ng et al., 2001). Subspaces can also be estimated beyond Euclidean data, for instance on symmetric positive definite (SPD) matrices datasets (Harandi et al., 2018). In all the previously cited examples, the search space—the space of all linear subspaces of dimension $q$ embedded in an ambient space of larger dimension $p$—is called the Grassmannian, or the Grassmann manifold (Bendokat et al., 2024) and is denoted $\operatorname{Gr}(p,q)$. Consequently, many machine learning methods can be recast as an optimization problem on Grassmann manifolds.111 Some alternative methods to subspace learning via Grassmannian optimization exist. Notably, a large family of methods—like partial least squares (PLS) (Geladi and Kowalski, 1986), independent component analysis (ICA) (Hyvärinen and Oja, 2000), canonical correlation analysis (CCA) (Hardoon et al., 2004) or projection pursuit (Huber, 1985)—builds a low-dimensional subspace out of the data sequentially. They first look for the best 1D approximation and then recursively add dimensions one-by-one via deflation schemes or orthogonality constraints (Lerman and Maunu, 2018b, Section II.C). Those sequential methods however suffer from many issues—from a subspace estimation point of view—due to their greedy nature. Those limitations are well described in Lerman and Maunu (2018b, Section III.B) and can be extended beyond robust subspace recovery. In light of the fundamental differences between those methods and the previously described subspace methods, we decide not to address sequential methods in this paper.
Learning a low-dimensional subspace from the data through Grassmannian optimization requires to choose the dimensionality $q$ a priori. This prerequisite has many important limitations, not only since a wrong choice of dimension might remove all the theoretical guarantees of the method (Lerman and Maunu, 2018b), but also because the observed dataset might simply not have a favored dimension, above which the added information is invaluable. [^2] Many generative models assume that the data lies on a lower dimensional subspace, up to isotropic Gaussian noise (Tipping and Bishop, 1999; Lerman et al., 2015). However, anisotropy is much more realistic (Maunu et al., 2019), which nuances the theoretical guarantees and removes hopes for a clear frontier between signal and noise at the intrinsic dimension. For these reasons, one might be tempted to “try” several dimensions—i.e. run the subspace optimization algorithm with different dimensions and choose the best one a posteriori, with cross-validation or statistical model selection techniques for instance. In addition to being potentially costly, such a heuristic raises an important issue that we later refer to as the nestedness issue: the optimal subspaces at different dimensions are not nested in general. This notably means that the data embeddings at two different dimensions might drastically differ, which is a pitfall for data analysis. For more details about the importance of nestedness in statistics, one can refer to Huckemann et al. (2010); Jung et al. (2012); Damon and Marron (2014); Huckemann and Eltzner (2018); Pennec (2018); Lerman and Maunu (2018b); Dryden et al. (2019); Yang and Vemuri (2021); Fan et al. (2022). We illustrate in Figure 1 the nestedness issue on toy datasets related to three important machine learning problems: robust subspace recovery, linear discriminant analysis and sparse spectral clustering (Lu et al., 2016; Wang et al., 2017). [^3] The methodology will be provided in due time.
Figure 1: Illustration of the subspace nestedness issue in three important machine learning problems: robust subspace recovery (left), linear discriminant analysis (middle) and sparse spectral clustering (right). For each dataset (top), we plot its projection onto the optimal 1D subspace (middle) and 2D subspace (bottom) obtained by solving the associated Grassmannian optimization problem. The 1D and 2D representations are inconsistent—in the sense that the 1D plot is not the projection of the 2D plot onto the horizontal axis—which is a pitfall for data analysis.
We can see that the scatter plots of the projected data for $q=1$ and $q=2$ are inconsistent—in that the 1D representation is not the projection of the 2D representation onto the horizontal axis.
In this paper, we propose a natural solution to the nestedness issue with a geometrical object that has not been much considered yet in machine learning: flags—i.e. sequences of nested subspaces of increasing dimension. The idea is to first select a sequence of candidate dimensions—the signature of the flag—then formulate a multilevel subspace learning criterion—with the flag trick—and finally optimize over the space of flags of the chosen signature, which has a manifold structure similar to the one of Grassmannians. The output subspaces are naturally nested, which solves the nestedness issue of subspace learning methods and provides the subspaces with a hierarchy that is good for interpretability (Huber, 1985, Section 7). Moreover, the nested representations of the data can be fit to general machine learning algorithms and combined via ensembling methods, in which case the weights can be interpreted as a measure of importance for the different dimensions, echoing the perspectives of the celebrated paper on the automatic choice of dimensionality in PCA (Minka, 2000). Beyond the change of paradigm (from subspaces to flags of subspaces), the main contribution of the paper is the flag trick: a generic method to convert a fixed-dimension subspace criterion into a multilevel flag criterion, without any nestedness issue.
The paper is organized as follows. In section 2, we provide some reminders on flag manifolds and describe a steepest descent optimization algorithm that we will use throughout this work. In section 3, we introduce the flag trick and propose an ensemble learning algorithm to leverage the hierarchical information of the flag into richer machine learning models. In section 4, we show the interest of the flag trick to several subspace learning methods such as robust subspace recovery, trace ratio and spectral clustering. In section 5, we conclude and discuss the limits and perspectives of such a framework. In appendix (Appendix A, Appendix B, Appendix C), we show how the flag trick can be used to develop two advanced optimization algorithms on flag manifolds (namely an iteratively reweighted least squares (IRLS) method for robust subspace recovery and a Newton-Lanczos method for trace ratio problems) that go beyond the limitations of simple algorithms like the steepest descent, and we provide the proofs that are not in the main body.
2 Reminders on flag manifolds
A flag is a sequence of nested linear subspaces of increasing dimension. This section introduces flag manifolds and provides the minimal tools to perform optimization on those spaces. Much more details and properties can be obtained in dedicated papers (Ye et al., 2022).
2.1 Flags in the scientific literature
By providing a natural parametrization for the eigenspaces of symmetric matrices, flags have long been geometrical objects of interest in the scientific community, with traditional applications in physics (Arnold, 1972) and linear algebra (Ammar and Martin, 1986). Modern computational methods on flag manifolds arose later with the seminal work of Edelman et al. (1998) who provided some optimization algorithms on matrix manifolds such as Grassmannians and Stiefel manifolds. Later, Nishimori et al. (2006) provided an optimization algorithm on flag manifolds for independent subspace analysis (Cardoso, 1998; Hyvärinen and Hoyer, 2000). Recently, flag manifolds played a central role in the investigation of a so-called curse of isotropy in principal component analysis (Szwagier and Pennec, 2024).
Flags as a sequence of nested subspaces were first introduced in the machine learning literature as robust prototypes to represent collections of subspaces of different dimensions (Draper et al., 2014; Mankovich et al., 2022). They were obtained via a sequential construction (where one dimension is added at a time), which can be problematic for greediness reasons (Huber, 1985; Lerman and Maunu, 2018b). Pennec (2018) was the first to show that PCA can be reformulated as an optimization problem on flag manifolds by summing the unexplained variance at different dimensions. This principle was recently applied to several variants of PCA under the name of flagification in Mankovich et al. (2024), showing an important robustness to outliers. Finally, some works compute distance-related quantities on flag-valued datasets, with notable applications in computer vision and hyperspectral imaging (Ma et al., 2021; Nguyen, 2022; Mankovich and Birdal, 2023; Szwagier and Pennec, 2023).
2.2 Definition and representation of flag manifolds
[…]
The set of flags of signature $(p,q_{1:d})$ is a smooth manifold (Ye et al., 2022), denoted here $\operatorname{Fl}(p,q_{1:d})$. Flag manifolds generalize Grassmannians—since $\leavevmode\nobreak\ {\operatorname{Fl}(p,(q,))=\operatorname{Gr}(p,q)}$—and therefore share many practical properties that are useful for optimization (Edelman et al., 1998). In the following, we will frequently use the following notations: $q_{0}:=0$, $q:=q_{d}$ and $q_{d+1}:=p$.
For computational and numerical reasons, flags are often represented as orthonormal $q$-frames. Those correspond to points on the Stiefel manifold $\operatorname{St}(p,q)=\{U\in\mathbb{R}^{p\times q}\colon U^{\top}U=I_{q}\}$). Let us define sequentially, for $k\in[1,d]$, $U_{k}\in\operatorname{St}(p,q_{k}-q_{k-1})$ such that $[U_{1}|\dots|U_{k}]$ is an orthonormal basis of $\mathcal{S}_{k}$ (this is possible thanks to the nestedness of the subspaces). Then, $U_{1:d}:=[U_{1}|\dots|U_{d}]\in\operatorname{St}(p,q)$ is a representative of the flag $\mathcal{S}_{1:d}$. We call it the Stiefel representation of flags. Such a representation is not unique—contrary to the projection representation defined previously—due to the rotational-invariance of orthonormal bases of subspaces. More precisely, if $U_{1:d}$ is a Stiefel representative of the flag $\mathcal{S}_{1:d}$, then for any set of orthogonal matrices $R_{k}\in\mathcal{O}(q_{k}-q_{k-1})$, the matrix $U^{\prime}_{1:d}:=[U_{1}R_{1}|\dots|U_{d}R_{d}]$ spans the same flag of subspaces $\mathcal{S}_{1:d}$. This provides flag manifolds with a quotient manifold structure (Edelman et al., 1998; Absil et al., 2009; Ye et al., 2022):
$\operatorname{Fl}(p,q_{1:d})\cong\operatorname{St}(p,q)\big{/}\left(\mathcal{O}(q_{1})\times\mathcal{O}(q_{2}-q_{1})\times\dots\times\mathcal{O}(q_{d}-q_{d-1})\right).$ (1)
Remark 1 (Orthogonal representation)
For computations, one might have to perform the orthogonal completion of some Stiefel representatives. Let $U_{1:d}:=[U_{1}|\dots|U_{d}]\in\operatorname{St}(p,q)$, then one denotes $U_{d+1}\in\operatorname{St}(p,p-q_{d})$ to be any orthonormal basis such that $U_{1:d+1}:=[U_{1}|\dots|U_{d}|U_{d+1}]\in\mathcal{O}(p)$. Such an orthogonal matrix $U_{1:d+1}$ will be called an orthogonal representative of the flag $\mathcal{S}_{1:d}$. In the following, we may abusively switch from one representation to the other since they represent the same flag.
2.3 Optimization on flag manifolds
There is a rich literature on optimization on smooth manifolds (Edelman et al., 1998; Absil et al., 2009; Boumal, 2023), and the particular case of flag manifolds has been notably addressed in Ye et al. (2022); Zhu and Shen (2024). Since flag manifolds can be represented as quotient spaces of Stiefel manifolds, which themselves are embedded in a Euclidean matrix space, one can develop some optimization algorithms without much difficulty. In this paper, we will use a steepest descent algorithm, which is drawn from several works (Chikuse, 2003; Nishimori et al., 2006; Ye et al., 2022; Zhu and Shen, 2024). Let $f\colon\operatorname{Fl}(p,q_{1:d})\to\mathbb{R}$ be a smooth function on a flag manifold, expressed in the Stiefel representation (e.g. $f(U_{1:d})=\sum_{k=1}^{d}\left\lVert{U_{k}}{U_{k}}^{\top}x\right\rVert$ for some $x\in\mathbb{R}^{p}$). Given $U_{1:d}\in\operatorname{St}(p,q)$, let $\operatorname{Grad}f(U_{1:d})=({\partial f}/{\partial U_{ij}})_{i,j=1}^{p,q}$ denote the (Euclidean) gradient of $f$. To “stay” on the manifold, one first computes the Riemannian gradient of $f$ at $U_{1:d}$, noted $\nabla f(U_{1:d})$. It can be thought of as a projection of the Euclidean gradient onto the tangent space and computed explicitly (Nishimori et al., 2006; Ye et al., 2022). Then, one moves in the opposite direction of $\nabla f(U_{1:d})$ with a so-called retraction, which is chosen to be the polar retraction of Zhu and Shen (2024, Eq. (49)), combined with a line-search. We iterate until convergence. The final steepest descent algorithm is described in Algorithm 1.
Algorithm 1 Steepest descent on flag manifolds
$f\colon\operatorname{Fl}(p,q_{1:d})\to\mathbb{R}$ a function, $U_{1:d}\in\operatorname{Fl}(p,q_{1:d})$ a flag (Stiefel representation)
for $t$ = 1, 2, … do
     $[G_{1}|\dots|G_{d}]\leftarrow\operatorname{Grad}f(U_{1:d})$triangleright Euclidean gradient
     $\nabla\leftarrow\bigl{[}G_{k}-\bigl{(}U_{k}{U_{k}}^{\top}G_{k}+\sum_{l\neq k}U_{l}{G_{l}}^{\top}U_{k}\bigr{)}\bigr{]}_{k=1\dots d}$triangleright Riemannian gradient
     $U_{1:d}\leftarrow\operatorname{polar}(U_{1:d+1}-\alpha\nabla)$triangleright polar retraction + line search
end for
$U_{1:d}^{*}\in\operatorname{Fl}(p,q_{1:d})$ an optimal flag
Remark 2 (Initialization)
We can initialize Algorithm 1 randomly via Chikuse (2003, Theorem 1.5.5), or choose a specific flag depending on the application, as we will see in section 4.
Remark 3 (Optimization variants)
Many extensions of Algorithm 1 can be considered: conjugate gradient (Ye et al., 2022), Riemannian trust region (Absil et al., 2009) etc. We can also replace the polar retraction with a geodesic step (Ye et al., 2022) or other retractions (Zhu and Shen, 2024).
3 The flag trick in theory
In this section, we motivate and introduce the flag trick to make subspace learning methods nested. The key result is Theorem 4, which states that the classical PCA at a fixed dimension can be converted into a nested multilevel method using nested projectors.
In the remaining of the paper, we assume that the data has been already centered around a point of interest (e.g. its mean or geometric median), so that we are only interested in fitting linear subspaces and not affine ones. One could directly include the center in the optimization variables—in which case the geometry would be the one of affine Grassmannians (Lim et al., 2019) or affine flags (Pennec, 2018)—but we don’t do it in this work for conciseness.
3.1 From subspaces to flags of subspaces: the seminal example of PCA
PCA is known as the eigendecomposition of the sample covariance matrix. Originally, it can be formulated as the search for a low dimensional subspace that minimizes the unexplained variance (or maximizes the explained variance). Let $X:=[x_{1}|\dots|x_{n}]\in\mathbb{R}^{p\times n}$ be a data matrix with $n$ samples, let $\mathcal{S}\in\operatorname{Gr}(p,q)$ be a $q$-dimensional subspace, and let $\Pi_{\mathcal{S}}\in\mathbb{R}^{p\times p}$ be the orthogonal projection matrix onto $\mathcal{S}$. Then PCA consists in the following optimization problem on Grassmannians:
$\mathcal{S}_{q}^{*}=\underset{\mathcal{S}\in\operatorname{Gr}(p,q)}{\operatorname{argmin}}\,\left\lVert X-\Pi_{\mathcal{S}}X\right\rVert_{F}^{2},$ (2)
where $\left\lVert M\right\rVert_{F}^{2}:=\operatorname{tr}\left(M^{\top}M\right)$ denotes the Frobenius norm. The solution to the optimization problem is the $q$-dimensional subspace spanned by the leading eigenvectors of the sample covariance matrix $S:=\frac{1}{n}XX^{\top}$, that we note $\mathcal{S}_{q}^{*}=\operatorname{Span}(v_{1},\dots,v_{q})$. It is unique when the sample eigenvalues $q$ and $q+1$ are distinct, which is almost sure when $q\leq\operatorname{rank}(S)$. We will assume to be in such a setting in the following for simplicity but it can be easily handled otherwise by “grouping” the repeated eigenvalues (cf. Szwagier and Pennec (2024, Theorem B.1)). In such a case, the principal subspaces are nested for increasing $q$, i.e., if $\mathcal{S}_{q}^{*}$ is the $q$-dimensional principal subspace, then for any $r>q$, one has $\mathcal{S}_{q}^{*}=\operatorname{Span}(v_{1},\dots,v_{q})\subset\operatorname{Span}(v_{1},\dots,v_{r})=\mathcal{S}_{r}^{*}$.
Another way of performing PCA is in a sequential manner (cf. footnote 1). We first estimate the 1D subspace $\mathcal{V}_{1}^{*}$ that minimizes the unexplained variance, then estimate the 1D subspace $\mathcal{V}_{2}^{*}$ that minimizes the unexplained variance while being orthogonal to the previous one, and so on and so forth. This gives the following constrained optimization problem on 1D Grassmannians:
$\mathcal{V}_{q}^{*}=\underset{\begin{subarray}{c}\mathcal{V}\in\operatorname{Gr}(p,1)\\ \mathcal{V}\perp\mathcal{V}_{q-1}\perp\dots\perp\mathcal{V}_{1}\end{subarray}}{\operatorname{argmin}}\,\left\lVert X-\Pi_{\mathcal{V}}X\right\rVert_{F}^{2}.$ (3)
This construction naturally yields a sequence of nested subspaces of increasing dimension—i.e. a flag of subspaces—best and best approximating the dataset:
$\{0\}\subset\mathcal{S}_{1}^{*}\subset\mathcal{S}_{2}^{*}\subset\dots\subset\mathcal{S}_{p-1}^{*}\subset\mathbb{R}^{p},\text{with }\mathcal{S}_{k}^{*}=\bigoplus_{l=1}^{k}\mathcal{V}_{l}^{*}.$ (4)
Those subspaces happen to be exactly the same as the ones obtained by solving the subspace learning optimization problem (2), although the way they are obtained (in a greedy manner) is different. This is generally not the case for other dimension reduction problems (for instance in robust subspace recovery, as it is raised in the final open questions of Lerman and Maunu (2018b)).
Hence, the optimal solution to the subspace learning formulation of PCA (2) is equivalent to the sequential formulation of PCA (3), and both yield a flag of subspaces best and best approximating the data. One can wonder if this result could be directly obtained by formulating an optimization problem on flag manifolds. The answer is yes, as first proven in Pennec (2018, Theorem 9) with an accumulated unexplained variance (AUV) technique, but there is not a unique way to do it. Motivated by the recent principal subspace analysis (Szwagier and Pennec, 2024), we propose in the following theorem a generic trick to formulate PCA as an optimization on flag manifolds.
Theorem 4 (Nested PCA with flag manifolds)
Let $X:=[x_{1}|\dots|x_{n}]\in\mathbb{R}^{p\times n}$ be a centered $p$-dimensional ($p\geq 2$) dataset with $n$ samples. Let $q_{1:d}:=(q_{1},q_{2},\cdots,q_{d})$ be a sequence of increasing dimensions such that $0<q_{1}<q_{2}<\cdots<q_{d}<p$. Let $S:=\frac{1}{n}XX^{\top}$ be the sample covariance matrix. Assume that it eigendecomposes as $S:=\sum_{j=1}^{p}\ell_{j}v_{j}{v_{j}}^{\top}$ where $\ell_{1}\geq\dots\geq\ell_{p}$ are the eigenvalues and $v_{1}\perp\dots\perp v_{p}$ are the associated eigenvectors. Then PCA can be reformulated as the following optimization problem on flag manifolds:
${\mathcal{S}_{1:d}^{*}}=\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\left\lVert X-\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}X\right\rVert_{F}^{2}.$ (5)
More precisely, one has ${\mathcal{S}_{1:d}^{*}}=\left(\operatorname{Span}(v_{1},\dots,v_{q_{1}}),\operatorname{Span}(v_{1},\dots,v_{q_{2}}),\dots,\operatorname{Span}(v_{1},\dots,v_{q_{d}})\right)$. The solution is unique if and only if $\ell_{q_{k}}\neq\ell_{q_{k+1}},\forall k\in[1,d]$.
Proof  One has:
$\left\lVert X-\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}X\right\rVert_{F}^{2}$ $=\operatorname{tr}\left(X^{\top}\left(I_{p}-\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}\right)^{2}X\right),$ (6)
$=\frac{1}{d^{2}}\operatorname{tr}\left(X^{\top}\left(\sum_{k=1}^{d}\left(I_{p}-\Pi_{\mathcal{S}_{k}}\right)\right)^{2}X\right),$ (7)
$=\frac{n}{d^{2}}\operatorname{tr}\left(W^{2}S\right),$ (8)
with $W=\sum_{k=1}^{d}(I_{p}-\Pi_{\mathcal{S}_{k}})$ and $S=\frac{1}{n}XX^{\top}$. Let $U_{1:d+1}:=[U_{1}|\dots|U_{d}|U_{d+1}]\in\mathcal{O}(p)$ be an orthogonal representative (cf. section 2) of the optimization variable $\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})$. Then one has $\leavevmode\nobreak\ {\Pi_{\mathcal{S}_{k}}=U_{1:d+1}\operatorname{diag}\left(I_{q_{k}},0_{p-q_{k}}\right){U_{1:d+1}}^{\top}}$. Therefore, one has $W=U_{1:d+1}\Lambda{U_{1:d+1}}^{\top}$, with $\leavevmode\nobreak\ {\Lambda=\operatorname{diag}\left(0\,I_{q_{1}},1\,I_{q_{2}-q_{1}},\dots,d\,I_{q_{d+1}-q_{d}}\right)}$. Hence, one has
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\left\lVert X-\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}X\right\rVert_{F}^{2}\Longleftrightarrow\underset{U\in\mathcal{O}(p)}{\operatorname{argmin}}\,\frac{n}{d^{2}}\operatorname{tr}\left(U\Lambda^{2}U^{\top}S\right).$ (9)
The latter problem is exactly the same as in Szwagier and Pennec (2024, Equation (19)), which solves maximum likelihood estimation for principal subspace analysis. Hence, one can conclude the proof on existence and uniqueness using Szwagier and Pennec (2024, Theorem B.1).
The key element of the proof of Theorem 4 is that averaging the nested projectors yields a hierarchical reweighting of the (mutually-orthogonal) principal subspaces. More precisely, the $k$-th principal subspace has weight $(k-1)^{2}$, and this reweighting enables to get a hierarchy of eigenspaces (Cunningham and Ghahramani, 2015; Pennec, 2018; Oftadeh et al., 2020). In the following, we note $\leavevmode\nobreak\ {\Pi_{\mathcal{S}_{1:d}}:=\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}}$ and call this symmetric matrix the average multilevel projector, which will be central in the extension of subspace methods into multilevel subspace methods.
3.2 The flag trick
As we will see in the following (section 4), many important machine learning problems can be formulated as the optimization of a certain function $f$ on Grassmannians. Theorem 4 shows that replacing the subspace projection matrix $\Pi_{\mathcal{S}}$ appearing in the objective function by the average multilevel projector $\Pi_{\mathcal{S}_{1:d}}$ yields a sequence of subspaces that meet the original objective of principal component analysis, while being nested. This leads us to introduce the flag trick for general subspace learning problems.
Definition 5 (Flag trick)
Let $p\geq 2$, $0<q<p$ and $q_{1:d}:=(q_{1},q_{2},\cdots,q_{d})$ be a sequence of increasing dimensions such that $0<q_{1}<q_{2}<\cdots<q_{d}<p$. The flag trick consists in replacing a subspace learning problem of the form:
$\underset{\mathcal{S}\in\operatorname{Gr}(p,q)}{\operatorname{argmin}}\,f(\Pi_{\mathcal{S}})$ (10)
with the following optimization problem:
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,f\left(\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}\right).$ (11)
Except for the very particular case of PCA (Theorem 4) where $f_{X}(\Pi)=\left\lVert X-\Pi X\right\rVert_{F}^{2}$, we cannot expect to have an analytic solution to the flag problem (11); indeed, in general, subspace problems do not even have a closed-form solution as we shall see in section 4. This justifies the introduction of optimization algorithms on flag manifolds like Algorithm 1.
Remark 6 (Flag trick vs. AUV)
The original idea of accumulated unexplained variance (Pennec, 2018) (and its subsequent application to several variants of PCA under the name of “flagification” (Mankovich et al., 2024)) consists in summing the subspace criteria at different dimensions, while the flag trick directly averages the orthogonal projection matrices that appear inside the objective function. While both ideas are equally worth experimenting with, we believe that the flag trick has a much wider reach. Indeed, from a technical viewpoint, the flag trick appears at the covariance level and directly yields a hierarchical reweighting of the principal subspaces. This reweighting is only indirect with the AUV—due to the linearity of the trace operator—and is not expected to happen for other methods than PCA. Notably, as we shall see in section 4, the flag trick enables to easily develop extensions of well-known methods involving PCA, like IRLS (Lerman and Maunu, 2018a) or Newton-Lanczos methods for trace ratio optimization (Ngo et al., 2012), and is closer in spirit to the statistical formulations of PCA (Szwagier and Pennec, 2024).
3.3 Multilevel machine learning
Subspace learning is often used as a preprocessing task before running a machine learning algorithm, notably to overcome the curse of dimensionality. One usually projects the data onto the optimal subspace $\mathcal{S}^{*}\in\operatorname{Gr}(p,q)$ and use the resulting lower-dimensional dataset as an input to a machine learning task like clustering, classification or regression (Bouveyron et al., 2019). Since the flag trick problem (11) does not output one subspace but a hierarchical sequence of nested subspaces, it is legitimate to wonder what to do with such a multilevel representation. In this subsection, we propose a general ensemble learning method to aggregate the hierarchical information coming from the flag of subspaces.
Let us consider a dataset $X:=[x_{1}|\dots|x_{n}]\in\mathbb{R}^{p\times n}$ (possibly with some associated labels $Y:=[y_{1}|\dots|y_{n}]\in\mathbb{R}^{m\times n}$). In machine learning, one often fits a model to the dataset by optimizing an objective function of the form $R_{X,Y}(g)=\frac{1}{n}\sum_{i=1}^{n}L(g(x_{i}),y_{i})$. With the flag trick, we get a filtration of projected data points $Z_{k}=\Pi_{\mathcal{S}_{k}^{*}}X,k\in[1,d]$ that can be given as an input to different machine learning algorithms. This yields optimal predictors $g_{k}^{*}=\operatorname{argmin}\,R_{Z_{k},Y}$ which can be aggregated via ensembling methods. For instance, voting methods choose the model with the highest performance on holdout data; this corresponds to selecting the optimal dimension $q^{*}\in q_{1:d}$ a posteriori, based on the machine learning objective. A more nuanced idea is the one of soft voting, which makes a weighted averaging of the predictions. The weights can be uniform, proportional to the performances of the individual models, or learned to maximize the performance of the weighted prediction (Perrone and Cooper, 1992). Soft voting gives different weights to the nested subspaces depending on their contribution to the ensembled prediction and therefore provides a soft measure of the relative importance of the different dimensions. In that sense, it goes beyond the classical manifold assumption stating that data has one intrinsic dimension, and instead proposes a soft blend between dimensions that is adapted to the learning objective. This sheds light on the celebrated paper of Minka for the automatic choice of dimensionality in PCA (Minka, 2000, Section 5). Many other ensembling methods are possible like gradient boosting, Bayesian model averaging and stacking. The whole methodology is summarized in Algorithm 2.
Algorithm 2 Flag trick combined with ensemble learning
$X:=[x_{1}|\dots|x_{n}]\in\mathbb{R}^{p\times n}$ a data matrix; $q_{1:d}:=(q_{1},\dots,q_{d})$ a flag signature; $f$ a subspace learning objective; (opt.) $Y:=[y_{1}|\dots|y_{n}]\in\mathbb{R}^{m\times n}$ a label matrix
${\mathcal{S}}_{1:d}^{*}\leftarrow\operatorname{argmin}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\,f(\Pi_{\mathcal{S}_{1:d}})$triangleright flag trick (11) + optimization (Alg. 1)
for k = 1 …d do
     $g_{k}^{*}\leftarrow\operatorname{fit}(\Pi_{\mathcal{S}_{k}^{*}}X,Y)$triangleright learning on $q_{k}$-dimensional projected data
     $Y_{k}^{*}\leftarrow g_{k}^{*}(\Pi_{\mathcal{S}_{k}^{*}}X)$triangleright prediction on $q_{k}$-dimensional projected data
end for
$Y^{*}\leftarrow\operatorname{ensembling}({Y}_{1}^{*},\dots,{Y}_{d}^{*})$triangleright weighted predictions
$Y^{*}$ the ensembled predictions
Algorithm 2 is a general proposition of multilevel machine learning with flags, but many other uses of the optimal flag $\mathcal{S}_{1:d}^{*}$ are possible, depending on the application. For instance, one may directly use the reweighted data matrix $\Pi_{\mathcal{S}_{1:d}^{*}}X$ as an input to the machine learning algorithm. This enables to fit only one model instead of $d$. One can also simply analyze the projected data qualitatively via scatter plots or reconstruction plots as evoked in section 1. The nestedness will automatically bring consistency contrarily to non-nested subspace methods, and therefore improve interpretability. Finally, many other ideas can be borrowed from the literature on subspace clustering and flag manifolds (Draper et al., 2014; Launay et al., 2021; Ma et al., 2021; Mankovich et al., 2022; Mankovich and Birdal, 2023; Mankovich et al., 2024), for instance the computation of distances between flags coming from different datasets as a multilevel measure of similarity between datasets.
4 The flag trick in action
In this section, we provide some applications of the flag trick to several learning problems. We choose to focus on subspace recovery, trace ratio and spectral clustering problems. Other ones, like domain adaptation, matrix completion and subspace tracking are developed or mentioned in the last subsection but not experimented for conciseness.
4.1 Outline and experimental setting
For each application, we first present the learning problem as an optimization on Grassmannians. Second, we formulate the associated flag learning problem by applying the flag trick (Definition 5). Third, we optimize the problem on flag manifolds with the steepest descent method (Algorithm 1)—more advanced algorithms are also derived in the appendix. Finally, we perform various nestedness and ensemble learning experiments via Algorithm 2 on both synthetic and real datasets.
The general methodology to compare Grassmann-based methods to flag-based methods is the following one. For each experiment, we first choose a flag signature $\leavevmode\nobreak\ {q_{1\rightarrow d}:=(q_{1},\dots,q_{d})}$, then we run independent optimization algorithms on $\operatorname{Gr}(p,q_{1}),\dots,\operatorname{Gr}(p,q_{d})$ (10) and finally we compare the optimal subspaces $\mathcal{S}_{k}^{*}\in\operatorname{Gr}(p,q_{k})$ to the optimal flag of subspaces $\mathcal{S}_{1:d}^{*}\in\operatorname{Fl}(p,q_{1\rightarrow d})$ obtained via the flag trick (11). To show the nestedness issue in Grassmann-based methods, we compute the subspace distances $\Theta(\mathcal{S}_{k}^{*},\mathcal{S}_{k+1}^{*})_{k=1\dots d-1}$, where $\Theta$ is the generalized Grassmann distance of Ye and Lim (2016, Eq. (14)). It consists in the $\ell_{2}$ norm of the principal angles, which can be obtained from the singular value decomposition (SVD) of the inner-products matrices ${U_{k}}^{\top}{U_{k+1}}$, where $U_{k}\in\operatorname{St}(p,q_{k})$ is an orthonormal basis of $\mathcal{S}_{k}^{*}$.
Regarding the implementation of the steepest descent algorithm on flag manifolds (Algorithm 1), we develop a new class of manifolds in PyManOpt (Boumal et al., 2014; Townsend et al., 2016), and run their SteepestDescent algorithm. Our implementation of the Flag class is based on the Stiefel representation of flag manifolds, detailed in section 2, with the retraction being the polar retraction. For the computation of the gradient, we use automatic differentiation with the autograd package. We could derive the gradients by hand from the expressions we get, but we use automatic differentiation as strongly suggested in PyManOpt’s documentation. Finally, the real datasets and the machine learning methods used in the experiments can be found in scikit-learn (Pedregosa et al., 2011).
4.2 The flag trick for robust subspace recovery problems
Let us consider a dataset that is a union of inliers and outliers—the inliers are assumed to lie near a low-dimensional subspace $\mathcal{S}$ while the outliers live in the ambient space. The aim of robust subspace recovery (RSR) is to recover $\mathcal{S}$ (Lerman and Maunu, 2018b). In that sense, RSR is an outlier-robust extension of classical dimension reduction methods like PCA. Without further specifications, the RSR problem might not be well-posed. Therefore, the works in this domain often have to make some assumptions on the inlier and outlier distributions in order to obtain some convergence and recovery guarantees. For instance, in Lerman and Maunu (2018b), it is assumed that the inliers “fill” the lower-dimensional subspace and that the outliers are not much “aligned”; this is rigorously defined in Lerman et al. (2015) and Maunu et al. (2019) through permeance and alignment statistics. A typical generative model following those assumptions is the Haystack model introduced in Lerman et al. (2015). The Haystack model assumes an isotropic Gaussian distribution on the subspace for the inliers and an isotropic Gaussian distribution on the (full) ambient space for the outliers. A more realistic model—the generalized Haystack model—is introduced in Maunu et al. (2019) to circumvent the simplistic nature of the Haystack model. This one assumes general (anisotropic) Gaussian distributions for the inliers and outliers. This makes the learning harder, since the anisotropy may keep the inliers from properly permeating the low-dimensional subspace—as discussed in section 1. Therefore, one has to make some stronger assumptions on the inlier-outlier ratio and the covariance eigenvalues distributions to derive some convergence and recovery guarantees.
Remark 7 (Parametrization of RSR generative models)
The inlier distribution in the Haystack model follows the isotropic PPCA model (Bouveyron et al., 2007, 2011), while it follows the PPCA model (Tipping and Bishop, 1999) in the case of the generalized Haystack model. Both models are a special case of the principal subspace analysis models (Szwagier and Pennec, 2024). However, as argued in Szwagier and Pennec (2024), while the Haystack model is parameterized with Grassmannians, the generalized Haystack model—which has more degrees of freedom accounting for the anisotropy—is parameterized with Stiefel manifolds. Therefore, from a statistical modeling perspective, it only makes sense to conduct subspace learning experiments on the Haystack model and not the generalized one.
4.2.1 Application of the flag trick to RSR
Among the large family of methods for robust subspace recovery (Lerman and Maunu, 2018b) we consider the one of least absolute deviation (LAD) minimization, which can be explicitly formulated as an optimization problem on Grassmannians. PCA minimizes the sum of squared Euclidean distances between the points and the subspace. In the case of an outlier-contaminated dataset, the squared Euclidean distances might penalize too much the outliers, and therefore make the optimal subspace too much influenced by the outliers. To circumvent this well-known sensitivity of squared norms to outliers, many works propose to minimize the sum of absolute Euclidean distances, which defines the LAD minimization problem:
$\underset{\mathcal{S}\in\operatorname{Gr}(p,q)}{\operatorname{argmin}}\,\sum_{i=1}^{n}\left\lVert x_{i}-\Pi_{\mathcal{S}}x_{i}\right\rVert_{2}.$ (12)
The latter has the interest of being rotationally invariant (Ding et al., 2006) but the drawback of being NP-hard (McCoy and Tropp, 2011; Lerman and Maunu, 2018b) and obviously non-convex since Grassmannians are not. A first body of works relaxes the problem, for instance by optimizing on the convex hull of Grassmannians (McCoy and Tropp, 2011; Xu et al., 2012; Zhang and Lerman, 2014; Lerman et al., 2015). A second body of works directly optimizes the LAD on Grassmannians, either with an IRLS algorithm (Lerman and Maunu, 2018a) or with a geodesic gradient descent (Maunu et al., 2019), both achieving very good results in terms of recovery and speed. The following proposition applies the flag trick to the LAD problem (12).
Proposition 8 (Flag trick for RSR)
The flag trick applied to LAD reads
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\sum_{i=1}^{n}\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2},$ (13)
and is equivalent to the following optimization problem:
$\underset{U_{1:d+1}\in\mathcal{O}(p)}{\operatorname{argmin}}\,\sum_{i=1}^{n}\sqrt{\sum_{k=1}^{d+1}\left(\frac{k-1}{d}\right)^{2}\left\lVert{U_{k}}^{\top}x_{i}\right\rVert_{2}^{2}}.$ (14)
Proof  The proof is given in Appendix (Appendix A).
Equation (14) tells us several things. First of all, due to the $((k-1)/d)^{2}$ weights, the points that are far from the center should be in the first principal subspaces. Second, the square root decreases the influence of the outlying points, compared to the nested PCA of Theorem 4. Third, although less obvious, we can see that numerical issues are less prone to happen with the flag trick (14) than with the LAD (12), since the quantity under the square root is zero only when the first subspace $\leavevmode\nobreak\ {\mathcal{S}_{1}=\operatorname{Span}(U_{1})}$ contains a data point. Therefore, whenever $q_{1}$ is smaller than the dimension $q$ that one would have tried for classical RSR, the non-differentiability and exploding-gradient issues caused by the square root are less likely. Finally, since the nested LAD minimization (13) is nothing but a robust version of the nested PCA of Theorem 4 ($\sum_{i=1}^{n}\|x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\|_{2}^{2}$), a natural idea can be to initialize the optimization algorithm with the nested PCA solution. This is what is done in Maunu et al. (2019) for LAD minimization, and it is coming with stronger recovery guarantees.
4.2.2 Nestedness experiments for RSR
[…]
Figure 2: Illustration of the nestedness issue in robust subspace recovery. Given a dataset consisting in a mixture of inliers (blue) and outliers (orange) we plot its projection onto the optimal 1D subspace and 2D subspace obtained by solving the associated Grassmannian optimization problem (12) or flag optimization problem (13). We can see that the Grassmann representations are not nested, while the flag representations are nested and robust to outliers.
Figure 3: Distributions of (increasing) Euclidean reconstruction errors $\|x_{i}-\Pi_{\mathcal{S}_{1:d}^{*}}x_{i}\|_{i=1\dots n}$ on the corrupted digits dataset for Grassmann and flag methods. While inliers and outliers are mixed with the subspace method, we can see a clear transition with flags. This can be explained by the multilevel nature of the flag trick.
We can see that the Grassmann-based projections are non-nested while their flag counterparts are not only nested but also robust to outliers. This could be explained by the nestedness constraint of flag manifolds which imposes the 2D subspace to contain the 1D subspace.
Second, we perform an outlier detection experiment. A common methodology to detect outliers in a corrupted dataset is to first look for an outlier-robust subspace and then plot the distribution of distances between the data points and their projection onto the recovered subspace. This distribution is expected to show a clear gap between the inliers and outliers (Vidal et al., 2016, Fig. 3.7). However, in practice, one does not know which subspace dimension $q$ to choose. If $q$ is too large, then the recovered subspace may contain both inliers and outliers, and therefore the distribution of distances might be roughly 00. In contrast, if $q$ is too small, then some inliers may lie too far from the recovered subspace and be detected as outliers. An idea in the spirit of the flag trick is to perform an average ensembling of the reconstruction errors. More specifically, if $\left\lVert x_{i}-\Pi_{\mathcal{S}}x_{i}\right\rVert_{2}$ is the classical measure of robust reconstruction error, then we compute $\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}$. Such a score extracts information from projections at different levels and might result in a richer multilevel analysis. We consider a dataset where the inliers are images of $8\times 8$ handwritten 00’s and outliers correspond to other digits from $1$ to $9$, all extracted from the classical handwritten digits dataset (Alpaydin and Kaynak, 1998). The ambient dimension is $p=64$, the number of inliers is $n_{\mathrm{in}}=90$ and the number of outliers is $n_{\mathrm{out}}=10$. The intrinsic dimensions that we try are $q_{1:3}=(1,2,5)$. We plot the reconstruction error for the points of the digits dataset on the optimal flag $\mathcal{S}_{1:3}^{*}\in\operatorname{Fl}(p,(1,2,5))$ in Figure 3. We compare it to the metric on $\mathcal{S}_{3}\in\operatorname{Gr}(p,q_{3})$. We can see that the flag trick enables to clearly distinguish the inliers from the outliers compared to the Grassmann-based method, which is a consequence of the multilevel nature of flags.
4.2.3 Discussion on RSR optimization and objective functions
An IRLS algorithm
In all the experiments of this paper, we use a steepest descent method on flag manifolds (Algorithm 1) to solve the flag problems. However, for the specific problem of RSR (13), we believe that more adapted algorithms should be derived, notably due to the non-differentiability and exploding-gradient issues caused by the square root. To that extent, we derive in appendix (Appendix A) an IRLS scheme (Algorithm 3) for RSR. In short, the RSR problem (13) can be reformulated as a weighted least squares problem $\sum_{i=1}^{n}w_{i}\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}^{2}$ with $w_{i}={1}/{\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}}$ and optimized iteratively, with explicit expressions obtained via our central Theorem 4. We insist on the fact that such an IRLS algorithm could not be developed with the flagification of Pennec (2018); Mankovich et al. (2024), since a sum of square roots does not correspond to a least-squares problem.
More RSR problems
In this work, we explore one specific problem of RSR for conciseness, but we could investigate many other related problems, including robust PCA. Notably, drawing from the Grassmann averages (GA) method (Hauberg et al., 2016), one could develop many new multilevel RSR and RPCA objective functions. The idea behind GA is to replace data points with 1D subspaces ($\mathcal{S}_{i}=\operatorname{Span}(x_{i})$) and then perform subspace averaging methods to find a robust prototype for the dataset. GA ends up solving problems of the form $\operatorname{argmin}_{\mathcal{S}\in\operatorname{Gr}(p,1)}\sum_{i=1}^{n}w_{i}\,\operatorname{dist}_{\operatorname{Gr}(p,1)}^{2}(\operatorname{Span}(x_{i}),\mathcal{S})$, where $w_{i}$ are some weights and $\operatorname{dist}_{\operatorname{Gr}(p,1)}$ is a particular subspace distance detailed in Hauberg et al. (2016). Using instead some multidimensional subspace distances, like the principal angles and its variants (Hamm and Lee, 2008; Ye and Lim, 2016), we can develop many variants of the Grassmann averages, of the form $\operatorname{argmin}_{\mathcal{S}_{1:d}\in\operatorname{Gr}(p,q)}\sum_{i=1}^{n}w_{i}\,\rho(\sqrt{{x_{i}}^{\top}\Pi_{\mathcal{S}}{x_{i}}})$, where $\rho\colon\mathbb{R}\to\mathbb{R}$ is a real function, like $\rho(x)=\arccos(x)$ if we want subspace-angle-like distances, $\rho(x)=-x^{2}$ if we want PCA-like solutions, and many other possible robust variants. Applying the flag trick to those problems yields the following robust multilevel problem: $\operatorname{argmin}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\sum_{i=1}^{n}w_{i}\rho(\sqrt{{x_{i}}^{\top}\Pi_{\mathcal{S}_{1:d}}{x_{i}}})$.
4.3 The flag trick for trace ratio problems
Trace ratio problems are ubiquitous in machine learning (Ngo et al., 2012). They write as:
$\underset{U\in\operatorname{St}(p,q)}{\operatorname{argmax}}\,\frac{\operatorname{tr}\left(U^{\top}AU\right)}{\operatorname{tr}\left(U^{\top}BU\right)},$ (15)
where $A,B\in\mathbb{R}^{p\times p}$ are positive semi-definite matrices, with $\operatorname{rank}(B)>p-q$.
A famous example of TR problem is Fisher’s linear discriminant analysis (LDA) (Fisher, 1936; Belhumeur et al., 1997). It is common in machine learning to project the data onto a low-dimensional subspace before fitting a classifier, in order to circumvent the curse of dimensionality. It is well known that performing an unsupervised dimension reduction method like PCA comes with the risks of mixing up the classes, since the directions of maximal variance are not necessarily the most discriminating ones (Chang, 1983). The goal of LDA is to use the knowledge of the data labels to learn a linear subspace that does not mix the classes. Let $\leavevmode\nobreak\ {X:=[x_{1}|\dots|x_{n}]\in\mathbb{R}^{p\times n}}$ be a dataset with labels $Y:=[y_{1}|\dots|y_{n}]\in{[1,C]}^{n}$. Let $\mu=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ be the dataset mean and $\mu_{c}=\frac{1}{\#\{i:y_{i}=c\}}\sum_{i:y_{i}=c}x_{i}$ be the class-wise means. The idea of LDA is to search for a subspace $\mathcal{S}\in\operatorname{Gr}(p,q)$ that simultaneously maximizes the projected between-class variance $\sum_{c=1}^{C}\|\Pi_{\mathcal{S}}\mu_{c}-\Pi_{\mathcal{S}}\mu\|_{2}^{2}$ and minimizes the projected within-class variance $\sum_{c=1}^{C}\sum_{i:y_{i}=c}\|\Pi_{\mathcal{S}}x_{i}-\Pi_{\mathcal{S}}\mu_{c}\|_{2}^{2}$. This can be reformulated as a trace ratio problem (15), with $A=\sum_{c=1}^{C}(\mu_{c}-\mu)(\mu_{c}-\mu)^{\top}$ and $B=\sum_{c=1}^{C}\sum_{i:y_{i}=c}(x_{i}-\mu_{c})(x_{i}-\mu_{c})^{\top}$.
More generally, a large family of dimension reduction methods can be reformulated as a TR problem. The seminal work of Yan et al. (2007) shows that many dimension reduction and manifold learning objective functions can be written as a trace ratio involving Laplacian matrices of attraction and repulsion graphs. Intuitively, those graphs determine which points should be close in the latent space and which ones should be far apart. Other methods involving a ratio of traces are multi-view learning (Wang et al., 2023), partial least squares (PLS) (Geladi and Kowalski, 1986; Barker and Rayens, 2003) and canonical correlation analysis (CCA) (Hardoon et al., 2004), although these methods are originally sequential problems (cf. footnote 1) and not subspace problems.
Classical Newton-like algorithms for solving the TR problem (15) come from the seminal works of Guo et al. (2003); Wang et al. (2007); Jia et al. (2009). The interest of optimizing a trace-ratio instead of a ratio-trace (of the form $\operatorname{tr}\left((U^{\top}BU)^{-1}(U^{\top}AU)\right)$), that enjoys an explicit solution given by a generalized eigenvalue decomposition, is also tackled in those papers. The repulsion Laplaceans (Kokiopoulou and Saad, 2009) instead propose to solve a regularized version $\operatorname{tr}\left(U^{\top}BU\right)-\rho\operatorname{tr}\left(U^{\top}AU\right)$, which enjoys a closed-form, but has a hyperparameter $\rho$, which is directly optimized in the classical Newton-like algorithms for trace ratio problems.
4.3.1 Application of the flag trick to trace ratio problems
The trace ratio problem (15) can be straightforwardly reformulated as an optimization problem on Grassmannians, due to the orthogonal invariance of the objective function:
$\underset{\mathcal{S}\in\operatorname{Gr}(p,q)}{\operatorname{argmax}}\,\frac{\operatorname{tr}\left(\Pi_{\mathcal{S}}A\right)}{\operatorname{tr}\left(\Pi_{\mathcal{S}}B\right)}.$ (16)
The following proposition applies the flag trick to the TR problem (16).
Proposition 9 (Flag trick for TR)
The flag trick applied to the TR problem (16) reads
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmax}}\,\frac{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}A\right)}{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}B\right)}.$ (17)
and is equivalent to the following optimization problem:
$\underset{U_{1:d}\in\operatorname{St}(p,q)}{\operatorname{argmax}}\,\frac{\sum_{k=1}^{d}(d-(k-1))\operatorname{tr}\left({U_{k}}^{\top}A{U_{k}}\right)}{\sum_{l=1}^{d}(d-(l-1))\operatorname{tr}\left({U_{l}}^{\top}B{U_{l}}\right)}.$ (18)
Proof  The proof is given in Appendix (Appendix B).
Equation (18) tells us several things. First, the subspaces $\leavevmode\nobreak\ {\operatorname{Span}(U_{1})\perp\dots\perp\operatorname{Span}(U_{d})}$ are weighted decreasingly, which means that they have less and less importance with respect to the TR objective. Second, we can see that the nested trace ratio problem (17) somewhat maximizes the numerator $\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}A\right)$ while minimizing the denominator $\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}B\right)$. Both subproblems have an explicit solution corresponding to our nested PCA Theorem 4. Hence, one can naturally initialize the steepest descent algorithm with the $q$ highest eigenvalues of $A$ or the $q$ lowest eigenvalues of $B$ depending on the application. For instance, for LDA, initializing Algorithm 1 with the highest eigenvalues of $A$ would spread the classes far apart, while initializing it with the lowest eigenvalues of $B$ would concentrate the classes, which seems less desirable since we do not want the classes to concentrate at the same point.
4.3.2 Nestedness experiments for trace ratio problems
For all the experiments of this subsection, we consider the particular TR problem of LDA, although many other applications (marginal Fisher analysis (Yan et al., 2007), local discriminant embedding (Chen et al., 2005) etc.) could be investigated similarly.
First, we consider a synthetic dataset with five clusters. The ambient dimension is $p=3$ and the intrinsic dimensions that we try are $q_{1:2}=(1,2)$. We adopt a preprocessing strategy similar to Ngo et al. (2012): we first center the data, then run a PCA to reduce the dimension to $n-C$ (if $n-C<p$), then construct the LDA scatter matrices $A$ and $B$, then add a diagonal covariance regularization of $10^{-5}$ times their trace and finally normalize them to have unit trace. We run Algorithm 1 on Grassmann manifolds to solve the TR maximization problem (16), successively for $q_{1}=1$ and $q_{2}=2$. Then we plot the projections of the data points onto the optimal subspaces. We compare them to the nested projections onto the optimal flag output by running Algorithm 1 on $\operatorname{Fl}(3,(1,2))$ to solve (17). The results are shown in Figure 4.
Figure 4: Illustration of the nestedness issue in linear discriminant analysis (trace ratio problem). Given a dataset with five clusters, we plot its projection onto the optimal 1D subspace and 2D subspace obtained by solving the associated Grassmannian optimization problem (16) or flag optimization problem (17). We can see that the Grassmann representations are not nested, while the flag representations are nested and well capture the distribution of clusters. In this example, it is less the nestedness than the rotation of the optimal axes inside the 2D subspace that is critical to the analysis of the Grassmann-based method.
Figure 5: Illustration of the nestedness issue in linear discriminant analysis (trace ratio problem) on the digits dataset. For $q_{k}\in q_{1:d}:=(1,2,\dots,63)$, we solve the Grassmannian optimization problem (16) on $\operatorname{Gr}(64,q_{k})$ and plot the subspace angles $\Theta(\mathcal{S}_{k}^{*},\mathcal{S}_{k+1}^{*})$ (left) and explained variances ${\operatorname{tr}(\Pi_{\mathcal{S}_{k}^{*}}XX^{\top})}/{\operatorname{tr}(XX^{\top})}$ (right) as a function of $k$. We compare those quantities to the ones obtained by solving the flag optimization problem (17). We can see that the Grassmann-based method is highly non-nested and even yields an extremely paradoxical non-increasing explained variance (cf. red circle on the right).
We can see that the Grassmann representations are non-nested while their flag counterparts perfectly capture the filtration of subspaces that best and best approximates the distribution while discriminating the classes. Even if the colors make us realize that the issue in this experiment for LDA is not much about the non-nestedness but rather about the rotation of the principal axes within the 2D subspace, we still have an important issue of consistency.
Second, we consider the (full) handwritten digits dataset (Alpaydin and Kaynak, 1998). It contains $8\times 8$ pixels images of handwritten digits, from 00 to $9$, almost uniformly class-balanced. One has $n=1797$, $p=64$ and $C=10$. We run a steepest descent algorithm to solve the trace ratio problem (17). We choose the full signature $q_{1:63}=(1,2,\dots,63)$ and compare the output flag to the individual subspaces output by running optimization on $\operatorname{Gr}(p,q_{k})$ for $q_{k}\in q_{1:d}$. We plot the subspace angles $\Theta(\mathcal{S}_{k}^{*},\mathcal{S}_{k+1}^{*})$ and the explained variance ${\operatorname{tr}(\Pi_{\mathcal{S}_{k}^{*}}XX^{\top})}/{\operatorname{tr}(XX^{\top})}$ as a function of the $k$. The results are illustrated in Figure 5. We see that the subspace angles are always positive and even very large sometimes with the LDA. Worst, the explained variance is not monotonous. This implies that we sometimes loose some information when increasing the dimension, which is extremely paradoxical.
Third, we perform some classification experiments on the optimal subspaces for different datasets. For different datasets, we run the optimization problems on $\operatorname{Fl}(p,q_{1:d})$, then project the data onto the different subspaces in $\mathcal{S}_{1:d}^{*}$ and run a nearest neighbors classifier with $5$ neighbors. The predictions are then ensembled (cf. Algorithm 2) by weighted averaging, either with uniform weights or with weights minimizing the average cross-entropy:
$w_{1}^{*},\dots,w_{d}^{*}=\underset{\begin{subarray}{c}w_{k}\geq 0\\ \sum_{k=1}^{d}w_{k}=1\end{subarray}}{\operatorname{argmin}}\,-\frac{1}{nC}\sum_{i=1}^{n}\sum_{c=1}^{C}y_{ic}\ln\left(\sum_{k=1}^{d}w_{k}y_{kic}^{*}\right),$ (19)
where $y_{kic}^{*}\in[0,1]$ is the predicted probability that $x_{i}\in\mathbb{R}^{p}$ belongs to class $c\in\{1\dots C\}$, by the classifier $g_{k}^{*}$ that is trained on $Z_{k}:={U_{k}^{*}}^{\top}X\in\mathbb{R}^{q_{k}\times n}$. One can show that the latter is a convex problem, which we optimize using the cvxpy Python package (Diamond and Boyd, 2016). We report the results in Table 1.
Table 1: Classification results for the TR problem on real datasets. For each method (Gr: Grassmann optimization (16), Fl: flag optimization (16), Fl-U: flag optimization + uniform soft voting, Fl-W: flag optimization + optimal soft voting (19)), we give the cross-entropy of the projected-predictions with respect to the true labels.
dataset $n$ $p$ $q_{1:d}$ Gr Fl Fl-U Fl-W weights
breast $569$ $30$ $(1,2,5)$ $0.0986$ $0.0978$ $0.0942$ $0.0915$ $(0.754,0,0.246)$
iris $150$ $4$ $(1,2,3)$ $0.0372$ $0.0441$ $0.0410$ $0.0368$ $(0.985,0,0.015)$
wine $178$ $13$ $(1,2,5)$ $0.0897$ $0.0800$ $0.1503$ $0.0677$ $(0,1,0)$
digits $1797$ $64$ $(1,2,5,10)$ $0.4507$ $0.4419$ $0.5645$ $0.4374$ $(0,0,0.239,0.761)$
The first example tells us that the optimal $5D$ subspace obtained by Grassmann optimization less discriminates the classes than the $5D$ subspace from the optimal flag. This may show that the flag takes into account some lower dimensional variability that enables to better discriminate the classes. We can also see that the uniform averaging of the predictors at different dimensions improves the classification. Finally, the optimal weights improve even more the classification and tell that the best discrimination happens by taking a soft blend of classifier at dimensions $1$ and $5$. Similar kinds of analyses can be made for the other examples.
4.3.3 Discussion on TR optimization and kernelization
A Newton algorithm
In all the experiments of this paper, we use a steepest descent method on flag manifolds (Algorithm 1) to solve the flag problems. However, for the specific problem of TR (17), we believe that more adapted algorithms should be derived to take into account the specific form of the objective function, which is classically solved via a Newton-Lanczos method (Ngo et al., 2012). To that extent, we develop in the appendix (Appendix B) an extension of the baseline Newton-Lanczos algorithm for the flag-tricked problem (17). In short, the latter can be reformulated as a penalized optimization problem of the form $\operatorname{argmax}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}}(A-\rho B)\right)}$, where $\rho$ is updated iteratively according to a Newton scheme. Once again, our central Theorem 4 enables to get explicit expressions for the iterations, which results without much difficulties in a Newton method, that is known to be much more efficient than first-order methods like the steepest descent.
A non-linearization via the kernel trick
The classical trace ratio problems look for linear embeddings of the data. However, in most cases, the data follow a nonlinear distribution, which may cause linear dimension reduction methods to fail. The kernel trick (Hofmann et al., 2008) is a well-known method to embed nonlinear data into a linear space and fit linear machine learning methods. As a consequence, we propose in appendix (Appendix B) a kernelization of the trace ratio problem (17) in the same fashion as the one of the seminal graph embedding work (Yan et al., 2007). This is expected to yield much better embedding and classification results.
4.4 The flag trick for spectral clustering problems
Spectral clustering (Ng et al., 2001) is a baseline clustering technique. It can be applied to general cluster distributions compared to the classical k-means (round clusters) and Gaussian mixture models (ellipsoidal clusters). Given a dataset $X:=[x_{1}|\dots|x_{n}]\in\mathbb{R}^{p\times n}$, spectral clustering roughly consists in the eigenvalue decomposition of a Laplacian matrix $L\in\operatorname{Sym}_{n}$ associated to a pairwise similarity matrix, for instance $M_{ij}=\exp(\left\lVert x_{i}-x_{j}\right\rVert^{2}/2\sigma^{2})$. The eigenvectors are then used as new embeddings for the data points, on which standard clustering algorithms like k-means can be performed. This method is closely related to the celebrated Laplacian eigenmaps (Belkin and Niyogi, 2003) which are used for nonlinear dimension reduction. The good performances of such a simple method as spectral clustering are theoretically justified by the particular structure of the Laplacian matrix $L$ in an ideal case—block-diagonal with a multiple eigenvalue related to the number of clusters (Ng et al., 2001). The recent sparse spectral clustering (SSC) method (Lu et al., 2016) builds on such an ideal case and encourages the block-diagonality by looking for a sparse and low-rank representation:
$\underset{U\in\operatorname{St}(p,q)}{\operatorname{argmin}}\,\langle UU^{\top},L\rangle_{F}+\beta\left\lVert UU^{\top}\right\rVert_{1},$ (20)
that they optimize over the convex hull of Grassmann manifolds with an ADMM algorithm.
4.4.1 Application of the flag trick to spectral clustering
The Grassmannian spectral clustering method (Wang et al., 2017) directly optimizes (20) on Grassmann manifolds:
$\underset{\mathcal{S}\in\operatorname{Gr}(p,q)}{\operatorname{argmin}}\,\langle\Pi_{\mathcal{S}},L\rangle_{F}+\beta\left\lVert\Pi_{\mathcal{S}}\right\rVert_{1}.$ (21)
The authors use a Riemannian trust region method (Absil et al., 2009) for optimization and show the interest of directly optimizing over the Grassmann manifold instead of convexifying the optimization space (Lu et al., 2016). The following proposition applies the flag trick to SSC.
Proposition 10 (Flag trick for SSC)
The flag trick applied to SSC (21) reads
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\langle\Pi_{\mathcal{S}_{1:d}},L\rangle_{F}+\beta\left\lVert\Pi_{\mathcal{S}_{1:d}}\right\rVert_{1}.$ (22)
and is equivalent to the following optimization problem:
$\underset{U_{1:d}\in\operatorname{St}(p,q)}{\operatorname{argmin}}\,\sum_{k=1}^{d}(d-(k-1))\operatorname{tr}\left({U_{k}}^{\top}LU_{k}\right)+\beta\left\lVert\sum_{k=1}^{d}(d-(k-1))U_{k}{U_{k}}^{\top}\right\rVert_{1}.$ (23)
Proof  The proof is given in Appendix (Appendix C).
We can see that the case $\beta=0$ corresponds to classical spectral clustering. Indeed, with a similar reasoning as in the proof of Theorem 4, we can easily show that the solution to problem (22) is explicit and corresponds to the flag of nested eigenspaces of $L$ (for increasing eigenvalues). Therefore, initializing the algorithm with the smallest $q$ eigenvectors of $L$ seems like a natural idea. Moreover, one may intuitively analyze the relative weighting of the mutually-orthogonal subspaces $\operatorname{Span}(U_{1})\perp\dots\perp\operatorname{Span}(U_{d})$ in (23) as favoring a model with $q_{1}$ clusters, and then adding successively $q_{k}-q_{k_{1}}$ clusters to improve the modeling of the Laplacian matrix, with a tradeoff between too much and too few clusters.
4.4.2 Nestedness experiments for spectral clustering
We consider a 3D extension of the classical two-moons dataset for clustering with $n=100$. The ambient dimension is $p=3$ and the intrinsic dimensions that we try are $q_{1:2}=(1,2)$. We adopt a pre-processing strategy similar to Lu et al. (2016); Wang et al. (2017), i.e. we compute the affinity matrix $W\in\mathbb{R}^{n\times n}$ using an exponential kernel with standard deviation being the median of the pairwise Euclidean distances between samples. Then we compute the normalized Laplacian matrix $L=I_{n}-D^{-\frac{1}{2}}LD^{-\frac{1}{2}}$ where $D\in\mathbb{R}^{n\times n}$ is a diagonal matrix with diagonal elements $D_{ii}=\sum_{j=1}^{n}w_{ij}$. We run Algorithm 1 on Grassmann manifolds to solve the sparse optimization problem (21), successively for $q_{1}=1$ and $q_{2}=2$. Then we plot the projections of the data points onto the optimal subspaces. We compare them to the nested projections onto the optimal flag output by running Algorithm 1 on $\operatorname{Fl}(3,(1,2))$ to solve (22). The results are shown in Figure 6.
Figure 6: Illustration of the nestedness issue in sparse spectral clustering. Given a dataset with two half-moon-shaped clusters, we plot its projection onto the optimal 1D subspace and 2D subspace obtained by solving the associated Grassmannian optimization problem (21) or flag optimization problem (22). We can see that the Grassmann representations are not nested, while the flag representations are nested and much better clustered. The last row corresponds to dividing the 2D embeddings by their norms, as commonly done in spectral clustering (Ng et al., 2001).
Figure 7: Illustration of the nestedness issue in sparse spectral clustering on the breast cancer dataset. For $q_{k}\in q_{1:d}:=(1,2,\dots,99)$, we solve the Grassmannian optimization problem (21) on $\operatorname{Gr}(100,q_{k})$ and plot the subspace angles $\Theta(\mathcal{S}_{k}^{*},\mathcal{S}_{k+1}^{*})$ (left) and explained variances ${\operatorname{tr}(\Pi_{\mathcal{S}_{k}^{*}}XX^{\top})}/{\operatorname{tr}(XX^{\top})}$ (right) as a function of $k$. We compare those quantities to the ones obtained by solving the flag optimization problem (22). We can see that the Grassmann-based method is highly non-nested while the flag of subspaces seems to be coherent with the cluster structure: the first axis is for the orange cluster while the second axis is for the blue cluster.
We can see that the Grassmann representations are not only non-nested, but also mix the two clusters in 2D, while the flag representations are nested and much better discriminate the clusters. This could be explained by the nestedness constraint of flag manifolds which imposes the 2D subspace to contain the 1D subspace.
Second, we consider the breast cancer dataset (Wolberg et al., 1995). It contains $569$ samples—from which we extract a subset of $100$ samples for computational time—with $30$ numerical attributes and two classes. One has $n=100$, $p=30$ and $C=2$. Then we run the steepest descent algorithm 1 to solve the SSC problem (22). We choose the full signature $q_{1:99}=(1,2,\dots,99)$ and compare the output flag to the individual subspaces output by running optimization on $\operatorname{Gr}(p,q_{k})$ for $q_{k}\in q_{1:d}$. We perform scatter plots in $1D$ and $2D$ and subspace error computations as a function of the $q_{k}$. The results are illustrated in Figure 7. The subspace angle plot tells us that the Grassmann spectral clustering yields highly non-nested subspaces, while the flags are by nature nested. The scatter plots show that the Grassmann representations are totally inconsistent, while the flag representations are consistent with the cluster structures.
Third, we perform some classification experiments on the optimal subspaces for different datasets. For different datasets, we run the optimization problems on $\operatorname{Fl}(p,q_{1:d})$, then project the data onto the different subspaces in $\mathcal{S}_{1:d}^{*}$, normalize the rows of their Stiefel representative $U_{k}^{*}\in\operatorname{St}(p,q_{k})$ to have unit Euclidean length and run a k-means algorithm with the number of clusters being equal to the number of classes. We report the clustering results in Table 2.
Table 2: Clustering results for the SSC problem on real datasets. For each method (Gr: Grassmann optimization (21), Fl: flag optimization (22)), we give the Rand index metric (Rand, 1971) of the projected-predictions with respect to the true cluster labels.
dataset $n$ $p$ $q_{1:d}$ Gr Fl
iris $99$ $4$ $(1,3,5)$ $0.8846$ $0.9371$
breast $100$ $30$ $(1,2,5)$ $0.5533$ $0.8181$
wine $99$ $13$ $(1,3,5)$ $0.6910$ $0.7108$
digits $100$ $64$ $(1,2,5,10)$ $0.9743$ $0.9820$
We can see that the flag approach yields more discriminating subspaces than the Grassmann approach. This can be explained by the multilevel nature of flags.
4.4.3 Discussion on SSC optimization
Similarly as in the previous examples, we can see that the steepest descent algorithm 1 might not be the most suited to solve the optimization problem (22), notably due to the $\ell_{1}$ penalization. Contrary to the RSR case in which we want to avoid the non-differentiable points (cf. subsubsection 4.2.3), here we would ideally like to attain those points, since they encode the desired sparsity of the Laplacian matrix. Therefore, we believe that more standard methods taking into account the $\ell_{1}$ norm to enforce the hierarchical sparsity of the flag of subspaces would be more adapted.
4.5 The flag trick for other machine learning problems
Subspace learning finds many applications beyond robust subspace recovery, trace ratio and spectral clustering problems, as evoked in section 1. The goal of this subsection is to provide a few more examples in brief, without experiments.
4.5.1 Domain adaptation
In machine learning, it is often assumed that the training and test datasets follow the same distribution. However, some domain shift issues—where training and test distributions are different—might arise, notably if the test data has been acquired from a different source (for instance a professional camera and a phone camera) or if the training data has been acquired a long time ago. Domain adaptation is an area of machine learning that deals with domain shifts, usually by matching the training and test distributions—often referred to as source and target distributions—before fitting a classical model (Farahani et al., 2021). A large body of works (called “subspace-based”) learn some intermediary subspaces between the source and target data, and perform the inference for the projected target data on these subspaces. The sampling geodesic flow (Gopalan et al., 2011) first performs a geodesic interpolation on Grassmannians between the source and target subspaces, then projects both datasets on (a discrete subset of) the interpolated subspaces, which results in a new representation of the data distributions, that can then be given as an input to a machine learning model. The higher the number of intermediary subspaces, the better the approximation, but the larger the dimension of the representation. The celebrated geodesic flow kernel (Boqing Gong et al., 2012) circumvents this issue by integrating the projected data onto the continuum of interpolated subspaces. This yields an inner product between infinite-dimensional embeddings that can be computed explicitly and incorporated in a kernel method for learning. The domain invariant projection (Baktashmotlagh et al., 2013) learns a domain-invariant subspace that minimizes the maximum mean discrepancy (MMD) (Gretton et al., 2012) between the projected source $X_{s}:=[x_{s1}|\dots|x_{sn_{s}}]\in\mathbb{R}^{p\times n_{s}}$ and target distributions $X_{t}:=[x_{t1}|\dots|x_{tn_{t}}]\in\mathbb{R}^{p\times n_{t}}$:
$\underset{U\in\operatorname{St}(p,q)}{\operatorname{argmin}}\,\operatorname{MMD}^{2}(U^{\top}X_{s},U^{\top}X_{t}),$ (24)
where
$\operatorname{MMD}(X,Y)=\left\lVert\frac{1}{n}\sum_{i=1}^{n}\phi(x_{i})-\frac{1}{m}\sum_{i=1}^{m}\phi(y_{i})\right\rVert_{\mathcal{H}}.$ (25)
This can be rewritten, using the Gaussian kernel function $\phi(x)\colon y\mapsto\exp\left(-\frac{x^{\top}y}{2\sigma^{2}}\right)$, as
$\underset{\mathcal{S}\in\operatorname{Gr}(p,q)}{\operatorname{argmin}}\,\frac{1}{n_{s}^{2}}\sum_{i,j=1}^{n_{s}}\exp\left(-\frac{(x_{si}-x_{sj})^{\top}\Pi_{\mathcal{S}}(x_{si}-x_{sj})}{2\sigma^{2}}\right)\\ +\frac{1}{n_{t}^{2}}\sum_{i,j=1}^{n_{t}}\exp\left(-\frac{(x_{ti}-x_{tj})^{\top}\Pi_{\mathcal{S}}(x_{ti}-x_{tj})}{2\sigma^{2}}\right)\\ -\frac{2}{n_{s}n_{t}}\sum_{i=1}^{n_{s}}\sum_{j=1}^{n_{t}}\exp\left(-\frac{(x_{si}-x_{tj})^{\top}\Pi_{\mathcal{S}}(x_{si}-x_{tj})}{2\sigma^{2}}\right).$ (26)
The flag trick applied to the domain invariant projection problem (26) yields:
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\frac{1}{n_{s}^{2}}\sum_{i,j=1}^{n_{s}}\exp\left(-\frac{(x_{si}-x_{sj})^{\top}\Pi_{\mathcal{S}_{1:d}}(x_{si}-x_{sj})}{2\sigma^{2}}\right)\\ +\frac{1}{n_{t}^{2}}\sum_{i,j=1}^{n_{t}}\exp\left(-\frac{(x_{ti}-x_{tj})^{\top}\Pi_{\mathcal{S}_{1:d}}(x_{ti}-x_{tj})}{2\sigma^{2}}\right)\\ -\frac{2}{n_{s}n_{t}}\sum_{i=1}^{n_{s}}\sum_{j=1}^{n_{t}}\exp\left(-\frac{(x_{si}-x_{tj})^{\top}\Pi_{\mathcal{S}_{1:d}}(x_{si}-x_{tj})}{2\sigma^{2}}\right),$ (27)
and can be rewritten as:
$\underset{U_{1:d}\in\operatorname{St}(p,q)}{\operatorname{argmin}}\,\frac{1}{{n_{s}}^{2}}\sum_{i,j=1}^{n_{s}}\exp\left(-\sum_{k=1}^{d}\frac{d+1-k}{d}\frac{\left\lVert{U_{k}}^{\top}(x_{si}-x_{sj})\right\rVert_{2}^{2}}{2\sigma^{2}}\right)\\ +\frac{1}{{n_{t}}^{2}}\sum_{i,j=1}^{n_{t}}\exp\left(-\sum_{k=1}^{d}\frac{d+1-k}{d}\frac{\left\lVert{U_{k}}^{\top}(x_{ti}-x_{tj})\right\rVert_{2}^{2}}{2\sigma^{2}}\right)\\ -\frac{2}{{n_{s}}{n_{t}}}\sum_{i=1}^{n_{s}}\sum_{j=1}^{n_{t}}\exp\left(-\sum_{k=1}^{d}\frac{d+1-k}{d}\frac{\left\lVert{U_{k}}^{\top}(x_{si}-x_{tj})\right\rVert_{2}^{2}}{2\sigma^{2}}\right).$ (28)
Some experiments similar to the ones of Baktashmotlagh et al. (2013) can be performed. For instance, one can consider the benchmark visual object recognition dataset of Saenko et al. (2010), learn nested domain invariant projections, fit some support vector machines to the projected source samples at increasing dimensions, and then perform soft-voting ensembling by learning the optimal weights on the target data according to Equation (19).
4.5.2 Low-rank decomposition
Many machine learning methods involve finding low-rank representations of a data matrix.
This is the case of matrix completion (Candès and Recht, 2012) problems where one looks for a low-rank representation of an incomplete data matrix by minimizing the discrepancy with the observed entries, and which finds many applications including the well-known Netflix problem. Although its most-known formulation is as a convex relaxation, it can also be formulated as an optimization problem on Grassmann manifolds (Keshavan et al., 2010; Boumal and Absil, 2011) to avoid optimizing the nuclear norm in the full space which can be of high dimension. The intuition is that a low-dimensional point can be described by the subspace it belongs to and its coordinates within this subspace. More precisely, the SVD-based low-rank factorization $M=UW$, with $M\in\mathbb{R}^{p\times n}$, $U\in\operatorname{St}(p,q)$ and $W\in\mathbb{R}^{q\times n}$ is orthogonally-invariant—in the sense that for any $R\in\mathcal{O}(q)$, one has $(UR)(R^{\top}W)=UW$. One could therefore apply the flag trick to such problems, with the intuition that we would try low-rank matrix decompositions at different dimensions. The application of the flag trick would however not be as straightforward as in the previous problems since the subspace-projection matrices $\Pi_{\mathcal{S}}:=UU^{\top}$ do not appear explicitly, and since the coefficient matrix $W$ also depends on the dimension $q$.
Many other low-rank problems can be formulated as a Grassmannian optimization. Robust PCA (Candès et al., 2011) looks for a low rank + sparse corruption factorization of a data matrix. Subspace Tracking (Balzano et al., 2010) incrementally updates a subspace from streaming and highly-incomplete observations via small steps on Grassmann manifolds.
4.5.3 Linear dimensionality reduction
Finally, many other general dimension reduction algorithms—referred to as linear dimensionality reduction methods (Cunningham and Ghahramani, 2015)—involve optimization on Grassmannians. For instance, linear dimensionality reduction encompasses the already-discussed PCA and LDA, but also many other problems like multi-dimensional scaling (Torgerson, 1952), slow feature analysis (Wiskott and Sejnowski, 2002), locality preserving projections (He and Niyogi, 2003) and factor analysis (Spearman, 1904).
5 Discussion
We introduced a general method to make subspace learning problems multilevel and nested. The nestedness issue of Grassmannian-based methods was clearly demonstrated on a variety of machine learning problems, yielding inconsistent scatter plots and paradoxical learning curves. The flag trick was shown to both solve the nestedness issue and provide solutions that are close from the original objective of the problems. As a positive side-effect, the flag trick sometimes learned substantially better subspaces than the original subspace problems, due to its multilevel nature which enables to aggregate information from several dimensions. When combined with ensemble learning, the flag trick showed improvement over subspace-based predictors that work at a fixed dimension, and raised interesting dimension-blending perspectives that challenge the manifold hypothesis. The code for reproducing the experiments can be found at https://github.com/tomszwagier/flag-trick.
One major limitation of this paper is the lack of in-depth study of the specific interest of the flag trick to each subspace learning method. We indeed focus on general practical results—the nestedness of flag-based methods with respect to classical subspace methods and the interest of ensembling dimensions instead of considering each dimension individually—and not on some application-specific questions, like the robustness of the flag to outliers for robust subspace recovery, the quality of the embedding for trace ratio problems, and the quality of the clustering for spectral clustering problems. We therefore strongly encourage the specialists to investigate with us the theoretical and practical interest of the flag trick for their subspace problems. We also encourage the Riemannian optimization community to develop with us new optimization algorithms on flag manifolds that are adapted to each subspace learning problem—for instance with robustness or sparsity-inducing properties. To that extent, we end this paper with a list of perspectives.
First, one should perform in-depth experiments to show the interest and the limitations of the flag trick for each specific subspace learning problem among—but not limited to—the list of examples developed in section 4, as well as their local, kernel and tangent space generalizations. Second, one should experiment with more complex and more efficient optimization algorithms on flag manifolds such as the ones developed in Appendix (IRLS in Appendix A and Newton in Appendix B) or the ones proposed in Ye et al. (2022); Zhu and Shen (2024) and develop new ones, like ADMM for sparse optimization. Third, we derived a very general principle for transitioning from fixed-dimension subspace methods to multi-dimensional flag methods, but this principle could be revisited specifically for each problem. This includes the problems that are not specifically formulated as a Grassmannian optimization, as long as they somewhat involve low-dimensional subspaces, like in domain adaptation (Gopalan et al., 2011; Boqing Gong et al., 2012) and sparse subspace clustering (Elhamifar and Vidal, 2013). Fourth, the question of how to choose the flag signature (i.e. the sequence of dimensions to try) remains open. A first step for PCA has been achieved with the principal subspace analysis (Szwagier and Pennec, 2024), where the dimensions are chosen based on the relative eigengaps of the covariance matrix. Other generative models should be designed specifically for each subspace learning problem in order to develop some appropriate guidelines to select the flag signature.
Acknowledgments and Disclosure of Funding
This work was supported by the ERC grant #786854 G-Statistics from the European Research Council under the European Union’s Horizon 2020 research and innovation program and by the French government through the 3IA Côte d’Azur Investments ANR-19-P3IA-0002 managed by the National Research Agency.
Appendix A Robust subspace recovery: extensions and proofs
A.1 An IRLS algorithm for robust subspace recovery
Iteratively reweighted least squares (IRLS) is a ubiquitous method to solve optimization problems involving $L^{p}$-norms. Motivated by the computation of the geometric median (Weiszfeld, 1937), IRLS is highly used to find robust maximum likelihood estimates of non-Gaussian probabilistic models (typically those containing outliers) and finds application in robust regression (Huber, 1964), sparse recovery (Daubechies et al., 2010) etc.
The recent fast median subspace (FMS) algorithm (Lerman and Maunu, 2018a), achieving state-of-the-art results in RSR uses an IRLS scheme to optimize the Least Absolute Deviation (LAD) (12). The idea is to first rewrite the LAD as
$\sum_{i=1}^{n}\left\lVert x_{i}-\Pi_{\mathcal{S}}x_{i}\right\rVert_{2}=\sum_{i=1}^{n}w_{i}(\mathcal{S})\left\lVert x_{i}-\Pi_{\mathcal{S}}x_{i}\right\rVert_{2}^{2},$ (29)
with $w_{i}(\mathcal{S})=\frac{1}{\left\lVert x_{i}-\Pi_{\mathcal{S}}x_{i}\right\rVert_{2}}$, and then successively compute the weights $w_{i}$ and update the subspace according to the weighted objective. More precisely, the FMS algorithm creates a sequence of subspaces $\mathcal{S}^{1},\dots,\mathcal{S}^{m}$ such that
$\mathcal{S}^{t+1}=\underset{\mathcal{S}\in\operatorname{Gr}(p,q)}{\operatorname{argmin}}\,\sum_{i=1}^{n}w_{i}(\mathcal{S}^{t})\left\lVert x_{i}-\Pi_{\mathcal{S}}x_{i}\right\rVert_{2}^{2}.$ (30)
This weighted least-squares problem enjoys a closed-form solution which relates to the eigenvalue decomposition of the weighted covariance matrix $\sum_{i=1}^{n}w_{i}(\mathcal{S}^{t})x_{i}{x_{i}}^{\top}$ (Vidal et al., 2016, Chapter 3.3).
We wish to derive an IRLS algorithm for the flag-tricked version of the LAD minimization problem (13). In order to stay close in mind to the recent work of Peng et al. (2023) who proved convergence of a general class of IRLS algorithms under some mild assumptions, we first rewrite (13) as
$\leavevmode\nobreak\ \underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\sum_{i=1}^{n}\rho(r(\mathcal{S}_{1:d},x_{i})),$ (31)
where $r(\mathcal{S}_{1:d},x)=\left\lVert x-\Pi_{\mathcal{S}_{1:d}}x\right\rVert_{2}$ is the residual and $\rho(r)=|r|$ is the outlier-robust loss function. Following Peng et al. (2023), the IRLS scheme associated with (31) is:
$\begin{cases}w_{i}^{t+1}=\rho^{\prime}(r(\mathcal{S}_{1:d}^{t},x_{i}))/r(\mathcal{S}_{1:d}^{t},x_{i})=1/\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2},\\ (\mathcal{S}_{1:d})^{t+1}=\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\sum_{i=1}^{n}w_{i}^{t+1}\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}^{2}.\end{cases}$ (32)
We now show that the second step enjoys a closed-form solution.
Theorem 11
The RLS problem
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\sum_{i=1}^{n}w_{i}\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}^{2}$ (33)
has a closed-form solution $\mathcal{S}_{1:d}^{*}\in\operatorname{Fl}(p,q_{1:d})$, which is given by the eigenvalue decomposition of the weighted sample covariance matrix $S_{w}=\sum_{i=1}^{n}w_{i}x_{i}{x_{i}}^{\top}=\sum_{j=1}^{p}\ell_{j}v_{j}{v_{j}}^{\top}$, i.e.
$\mathcal{S}_{k}^{*}=\operatorname{Span}(v_{1},\dots,v_{q_{k}})\quad(k=1\mathinner{\ldotp\ldotp}d).$ (34)
Proof  One has
$\sum_{i=1}^{n}w_{i}\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}^{2}=\operatorname{tr}\left((I-\Pi_{\mathcal{S}_{1:d}})^{2}\left(\sum_{i=1}^{n}w_{i}x_{i}{x_{i}}^{\top}\right)\right).$ (35)
Therefore, we are exactly in the same case as in Theorem 4, if we replace $XX^{\top}$ with the reweighted covariance matrix $\sum_{i=1}^{n}w_{i}x_{i}{x_{i}}^{\top}$. This does not change the result, so we conclude with the end of the proof of Theorem 4 (which itself relies on Szwagier and Pennec (2024)).
Hence, one gets an IRLS scheme for the LAD minimization problem. One can modify the robust loss function $\rho(r)=|r|$ by a Huber-like loss function to avoid weight explosion. Indeed, one can show that the weight $w_{i}:=1/\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}$ goes to infinity when the first subspace $\mathcal{S}_{1}$ of the flag gets close to $x_{i}$ Therefore in practice, we take
$\rho(r)=\begin{cases}r^{2}/(2p\delta)&\text{if }|r|<=p\delta,\\ r-p\delta/2&\text{if }|r|>p\delta.\end{cases}$ (36)
This yields
$w_{i}=1/\max\left(p\delta,1/\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}\right).$ (37)
The final proposed scheme is given in Algorithm 3, named fast median flag (FMF), in reference to the fast median subspace algorithm of Lerman and Maunu (2018a).
Algorithm 3 Fast median flag
$X\in\mathbb{R}^{p\times n}$ (data), $\quad q_{1}<\dots<q_{d}$ (signature), $\quad t_{max}$ (max number of iterations), $\quad\eta$ (convergence threshold), $\quad\varepsilon$ (Huber-like saturation parameter)
$U\in\operatorname{St}(p,q)$
$t\leftarrow 0,\quad\Delta\leftarrow\infty,\quad U^{0}\leftarrow\operatorname{SVD}(X,q)$
while $\Delta>\eta$ and $t<t_{max}$ do
     $t\leftarrow t+1$
     $r_{i}\leftarrow\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}$
     $y_{i}\leftarrow{x_{i}}/{\max(\sqrt{r_{i}},\varepsilon)}$
     $U^{t}\leftarrow\operatorname{SVD(Y,q)}$
     $\Delta\leftarrow\sqrt{\sum_{k=1}^{d}\Theta(U^{t}_{q_{k}},U^{t-1}_{q_{k}})^{2}}$
end while
We can easily check that FMF is a direct generalization of FMS for Grassmannians (i.e. when $d=1$).
Remark 12
This is far beyond the scope of the paper, but we believe that the convergence result of Peng et al. (2023, Theorem 1) could be generalized to the FMF algorithm, due to the compactness of flag manifolds and the expression of the residual function $r$.
A.2 Proof of Proposition 8
Let $\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})$ and $U_{1:d+1}:=[U_{1}|U_{2}|\dots|U_{d}|U_{d+1}]\in\mathcal{O}(p)$ be an orthogonal representative of $\mathcal{S}_{1:d}$. One has:
$\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}$ $=\sqrt{{(x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i})}^{\top}(x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i})},$ (38)
$=\sqrt{{x_{i}}^{\top}{(I_{p}-\Pi_{\mathcal{S}_{1:d}})}^{2}x_{i}},$ (39)
$=\sqrt{{x_{i}}^{\top}{\left(I_{p}-\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}\right)}^{2}x_{i}},$ (40)
$=\sqrt{\frac{1}{d^{2}}{x_{i}}^{\top}{\left(\sum_{k=1}^{d}(I_{p}-\Pi_{\mathcal{S}_{k}})\right)}^{2}x_{i}},$ (41)
$=\sqrt{\frac{1}{d^{2}}{x_{i}}^{\top}\left(\sum_{k=1}^{d+1}(k-1)U_{k}{U_{k}}^{\top}\right)^{2}x_{i}},$ (42)
$=\sqrt{\frac{1}{d^{2}}{x_{i}}^{\top}\left(\sum_{k=1}^{d+1}(k-1)^{2}U_{k}{U_{k}}^{\top}\right)x_{i}},$ (43)
$=\sqrt{\sum_{k=1}^{d+1}\left(\frac{k-1}{d}\right)^{2}{x_{i}}^{\top}\left(U_{k}{U_{k}}^{\top}\right)x_{i}},$ (44)
$\left\lVert x_{i}-\Pi_{\mathcal{S}_{1:d}}x_{i}\right\rVert_{2}$ $=\sqrt{\sum_{k=1}^{d+1}\left(\frac{k-1}{d}\right)^{2}\left\lVert{U_{k}}^{\top}x_{i}\right\rVert_{2}^{2}},$ (45)
which concludes the proof.
Appendix B Trace Ratio problems: extensions and proofs
B.1 A Newton method for multilevel trace ratio problems
We wish to solve the following optimization problem (cf. Equation 17):
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmax}}\,\frac{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}A\right)}{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}B\right)}.$ (46)
We draw from Ngo et al. (2012) (itself based on Wang et al. (2007); Jia et al. (2009)) to develop an efficient algorithm for solving this problem.
Let $\phi\colon\operatorname{Fl}(p,q_{1:d})\ni\mathcal{S}_{1:d}\mapsto{\operatorname{tr}\left(\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}A\right)}/{\operatorname{tr}\left(\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}B\right)}$. One can see that $\phi$ is well defined on $\operatorname{Fl}(p,q_{1:d})$ if and only if $\operatorname{rank}(B)>p-q_{d}$, which we assume in the following. $\phi$ is continuous on $\operatorname{Fl}(p,q_{1:d})$ which is a smooth compact manifold, therefore it admits a maximum $\rho^{*}$. As a consequence, for any ${\mathcal{S}_{1:d}}\in\operatorname{Fl}(p,q_{1:d})$, one has $\phi(\mathcal{S}_{1:d})\leq\rho^{*}$, which is equivalent to $\leavevmode\nobreak\ {\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}}(A-\rho^{*}B)\right)\leq 0}$, and since the maximum is attained, lets say at $\mathcal{S}_{1:d}^{*}:=(\mathcal{S}_{1}^{*},\dots,\mathcal{S}_{d}^{*})$, one has ${\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}^{*}}(A-\rho^{*}B)\right)=0}$.
Using a similar reasoning as in the proof of Theorem 4 (itself based on Szwagier and Pennec (2024)), we can show that for any $\rho\in\mathbb{R}$, the function $\mathcal{S}_{1:d}\mapsto\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}}(A-\rho B)\right)$ is maximized in ${\mathcal{S}_{1:d}^{*}}=\left(\operatorname{Span}(v_{1},\dots,v_{q_{1}}),\operatorname{Span}(v_{1},\dots,v_{q_{2}}),\dots,\operatorname{Span}(v_{1},\dots,v_{q_{d}})\right)$, where $v_{1},\dots,v_{p}$ are the eigenvectors associated with the ordered-descending eigenvalues $\ell_{1},\dots,\ell_{p}$. The maximum is simply given by $f(\rho):=\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}^{*}}(A-\rho B)\right)=\sum_{k=1}^{d}\sum_{j=1}^{q_{k}}\ell_{j}$ and it verifies $f(\rho^{*})=0$.
Let us now show that $\rho^{*}$ is the unique root of $f$. Let us introduce $\ell_{j}(\cdot)$, the function that maps a symmetric matrix to its $j$-th largest eigenvalue, and similarly $\mathcal{S}_{k}^{*}(\cdot)$, the function that maps a symmetric matrix to the subspace spanned by the eigenvectors associated with the $q_{k}$ largest eigenvalues. Then one can show that the derivative of $\sum_{j=1}^{q_{k}}\ell_{j}(A-\rho B)$ with respect to $\rho$ is $-\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}^{*}(A-\rho B)}B\right)$ (Ngo et al., 2012). Therefore, the derivative of $f$ is $f^{\prime}(\rho)=-\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}^{*}(A-\rho B)}B\right)$, which is everywhere strictly negative since $B$ is positive semidefinite and $\operatorname{rank}(B)>p-q_{d}$. Therefore, f is a strictly decreasing function. Hence, $\rho^{*}$ is the unique root of $f$.
Now, the idea is to develop a Newton method to find the root $\rho^{*}$. The optimal flag will then be given by the eigenvalue decomposition of $A-\rho^{*}B$. One has the following iteration:
$\rho^{t+1}=\rho^{t}-\frac{f(\rho^{t})}{f^{\prime}(\rho^{t})}=\frac{\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}(A-\rho^{t}B)}A\right)}{\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}(A-\rho^{t}B)}B\right)}$ (47)
Finally, in short, one can use similar considerations as in Ngo et al. (2012) to initialize the algorithm. If we assume that $B$ is positive definite, then we can find $Z\in\operatorname{GL}_{p}$ such that $Z^{\top}BZ=I_{p}$ and $Z^{\top}AZ=\Lambda$, where $\Lambda:=\operatorname{diag}\left(\lambda_{1}\geq\dots\geq\lambda_{p}\right)$ are the (ordered descending) generalized eigenvalues of the pencil $(A,B)$. Then, using Sylvester’s law of inertia, one gets that the number of negative eigenvalues of $Z^{\top}(A-\rho B)Z$ and of $\Lambda-\rho I_{p}$ are the same. Therefore, for $\rho=\lambda_{1}$, all the eigenvalues of $A-\rho B$ are nonpositive, so $f(\rho)=\sum_{k=1}^{d}\sum_{j=1}^{q_{k}}\lambda_{j}(A-\rho B)\leq 0$. Similarly, for $\rho=\lambda_{q_{d}}$, the $q_{d}$ largest eigenvalues of $A-\rho B$ are nonnegative, so in particular for all $k\in[1,d]$, the $q_{k}$ largest eigenvalues of $A-\rho B$ are nonnegative, so in the end $f(\rho)\geq 0$. To conclude, if $B$ is positive definite, then one has $\rho^{*}\in[\lambda_{q_{d}}(A,B),\lambda_{1}(A,B)]$. This can be used as an initialization to the Newton method, although in practice a random initialization is performed in practice in Ngo et al. (2012, Algorithm 4.1).
The final proposed scheme is given in Algorithm 4, called flag iterative trace ratio in reference to the iterative trace ratio of Wang et al. (2007).
Algorithm 4 Flag iterative trace ratio
$A,B\in\operatorname{SPD}_{p}$ (can find tighter restrictions), $\quad q_{1}<\dots<q_{d}$ (signature), $\quad\eta$ (convergence threshold)
$U\in\operatorname{St}(p,q)$
$t\leftarrow 0,\Delta\leftarrow\infty,\rho^{0}\in\left[\lambda_{q}(A,B),\lambda_{1}(A,B)\right]$
while $\Delta>\eta$ do
     $U^{t},\Lambda^{t}\leftarrow\operatorname{EVD}(A-\rho^{t}B,q)$
     $\rho^{t}\leftarrow\frac{\sum_{k=1}^{d}\operatorname{tr}\left({U^{t}_{q_{k}}}^{\top}AU^{t}_{q_{k}}\right)}{\sum_{k=1}^{d}\operatorname{tr}\left({U^{t}_{q_{k}}}^{\top}BU^{t}_{q_{k}}\right)}$
     $\Delta\leftarrow\sqrt{\sum_{k=1}^{d}\Theta(U^{t}_{q_{k}},U^{t-1}_{q_{k}})^{2}}$
end while
B.2 Kernelization (LDA and Graph Embedding)
Similarly to Yan et al. (2007); Wang et al. (2007), if the matrices $A$ and $B$ can be rewritten as $XL_{A}X^{\top}$ and $XL_{B}X^{\top}$ where $L_{A},L_{B}\in\operatorname{Sym}_{n}$ are Laplacian matrices of some connectivity graphs associated to the dataset, then one can apply the kernel trick to non-linearly map the data to a high-dimensional Hilbert space $(\mathcal{H},\langle\cdot,\cdot\rangle$. Let $\mathcal{K}$ be the chosen kernel, like for instance the RBF $\mathcal{K}(x,y)=e^{-\gamma\|x-y\|^{2}}$. Let ${K}:=(K(x_{i},x_{j}))_{i,j\in[1,n]}\in\operatorname{Sym}_{n}$. Then, by noting $\phi:\mathbb{R}^{p}\to\mathcal{H}$ the nonlinear map $\phi(x)=\mathcal{K}_{x}=\mathcal{K}(x,\cdot)$, one has $\langle\phi(x),\phi(y)\rangle=\mathcal{K}(x,y)$.
Let $K:=Q\Lambda Q^{\top}$ be an eigendecomposition of $K$, with $Q\in\mathcal{O}(n)$ and $\Lambda\in\operatorname{diag}(\mathbb{R}^{n})$. Then the trace ratio problem in the RKHS can be rewritten as:
$\underset{{\mathcal{S}_{1:d}}\in\operatorname{Flag}(n,q_{1:d})}{\operatorname{argmax}}\,\frac{\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}}A\right)}{\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}}B\right)},$ (48)
with
$A$ $=\Lambda^{1/2}Q^{\top}L_{A}Q\Lambda^{1/2}$
$B$ $=\Lambda^{1/2}Q^{\top}L_{B}Q\Lambda^{1/2}$
Then, noting $\Gamma:=[\Gamma_{1}|\dots|\Gamma_{d}]=[U_{1}|\dots|U_{d}]$, one get for any $x\in\mathbb{R}^{p},\Pi^{*}(x)=\sum_{i}\Gamma_{\dot{i}}K(x_{i},x)$, and therefore $\Pi^{*}(X)=\Gamma^{\top}K\in\mathbb{R}^{q\times n}$ and for any new data points, $x^{\prime}_{1},\dots,x^{\prime}_{m}\in\mathbb{R}^{p}$, writing ${K}^{\prime}=(K(x^{\prime}_{i},x_{j}))_{i\in[1,m],j\in[1,n]}\in\mathbb{R}^{m\times n}$, one has ${\Pi^{*}}(X^{\prime})=\Gamma^{\top}{K^{\prime}}^{\top}\in\mathbb{R}^{q\times m}$.
B.3 Proof of Proposition 9
Let $\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})$ and $U_{1:d+1}:=[U_{1}|U_{2}|\dots|U_{d}|U_{d+1}]\in\mathcal{O}(p)$ be an orthogonal representative of $\mathcal{S}_{1:d}$. One has:
$\frac{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}A\right)}{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}B\right)}$ $=\frac{\operatorname{tr}\left(\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}A\right)}{\operatorname{tr}\left(\frac{1}{d}\sum_{l=1}^{d}\Pi_{\mathcal{S}_{l}}B\right)},$ (49)
$=\frac{\operatorname{tr}\left(\sum_{k=1}^{d+1}(d-(k-1))U_{k}{U_{k}}^{\top}A\right)}{\operatorname{tr}\left(\sum_{l=1}^{d+1}(d-(l-1))U_{l}{U_{l}}^{\top}B\right)},$ (50)
$\frac{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}A\right)}{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}B\right)}$ $=\frac{\sum_{k=1}^{d+1}(d-(k-1))\operatorname{tr}\left(U_{k}{U_{k}}^{\top}A\right)}{\sum_{l=1}^{d+1}(d-(l-1))\operatorname{tr}\left(U_{l}{U_{l}}^{\top}B\right)},$ (51)
which concludes the proof.
Appendix C Spectral clustering: extensions and proofs
C.1 Proof of Proposition 10
Let $\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})$ and $U_{1:d+1}:=[U_{1}|U_{2}|\dots|U_{d}|U_{d+1}]\in\mathcal{O}(p)$ be an orthogonal representative of $\mathcal{S}_{1:d}$. One has:
$\langle\Pi_{\mathcal{S}_{1:d}},L\rangle_{F}+\beta\left\lVert\Pi_{\mathcal{S}_{1:d}}\right\rVert_{1}$ $=\left\langle\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}},L\right\rangle_{F}+\beta\left\lVert\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}\right\rVert_{1},$ (52)
$=\frac{1}{d}\left(\left\langle\sum_{k=1}^{d+1}(d-(k-1))U_{k}{U_{k}}^{\top},L\right\rangle_{F}+\beta\left\lVert\sum_{k=1}^{d+1}(d-(k-1))U_{k}{U_{k}}^{\top}\right\rVert_{1}\right),$ (53)
$=\frac{1}{d}\left(\sum_{k=1}^{d+1}(d-(k-1))\left\langle U_{k}{U_{k}}^{\top},L\right\rangle_{F}+\beta\left\lVert\sum_{k=1}^{d+1}(d-(k-1))U_{k}{U_{k}}^{\top}\right\rVert_{1}\right),$ (54)
$\langle\Pi_{\mathcal{S}_{1:d}},L\rangle_{F}+\beta\left\lVert\Pi_{\mathcal{S}_{1:d}}\right\rVert_{1}$ $=\frac{1}{d}\left(\sum_{k=1}^{d+1}(d-(k-1))\operatorname{tr}\left({U_{k}}^{\top}LU_{k}\right)+\beta\left\lVert\sum_{k=1}^{d+1}(d-(k-1))U_{k}{U_{k}}^{\top}\right\rVert_{1}\right),$ (55)
which concludes the proof.